Iodine converts from the molecular form to the atomic form as follows:

I2 (g) ⇌ 2I (g) at 500 K.

The initial [I2] = 0.45 M and initial [I] = 0.

What is [I] at equilibrium? Kc = 5.6 x 10-12 at 500K.

Hint:

You will need to set up an ICE chart and solve a quadratic equation for this one.
Answer #1

. = a1 kci s-661012 initial charge Iz Dusan -ze tare eab 0.4S-re Kc = S-681012 - = LI2) Oulus-re a 426² = 5-6x10le (6.45-) so